The air temperature is 20 oC. A stationary observer sees a police car receding into the distance with a speed of 30 m/s with it’s siren blaring. If the police siren is designed to produce a sound of frequency 900 Hz, calculate the frequency of the sound heard by the stationary observer.


1
Expert's answer
2021-01-27T16:34:21-0500

"v=343.1\ (m/s)\ \ (t=20°C)"


"\nu=\nu_0\cdot\frac{v}{v+u}=900\cdot\frac{343.1}{343.1+30}=827.6(Hz)" .Answer